OpenStudy (clamin):

Find the value of c that completes the square. y^2-34y+c

OpenStudy (anonymous):

ay^2 + by + c... the value of c needed to complete the square is: \[c=\frac{ b }{ ? }\]

OpenStudy (anonymous):

\[c=\frac{ b }{ 2a }\]

TheSmartOne (thesmartone):

c=(-34/2(1))^2 c=-17^2 c=289

TheSmartOne (thesmartone):

So your final answer will be y^2-34y+289

TheSmartOne (thesmartone):

If you factored it it would be (y-17)^2